### v1

@@ Abstract
GPU-enhanced architectures are now dominant in HPC systems, but message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches that lower such costs. We compare and contrast stream/graph- and kernel-triggered MPI communication abstractions, whose principal purpose is to enhance the performance of communication when GPU kernels create or consume data for transfer through MPI operations. Researchers and practitioners have proposed multiple potential APIs for stream and/or kernel triggering that span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, and explicit MPI stream/queue objects. Designs breaking backward compatibility with MPI are duly noted. Some of these strengthen or weaken the semantics of MPI operations. A key contribution of this paper is to promote community convergence toward a stream- and/or kernel-triggering abstraction by highlighting the common and differing goals and contributions of existing abstractions. We describe the design space in which these abstractions reside, their implicit or explicit use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discuss their potential for added performance, how usable these abstractions are, and where functional and/or semantic gaps exist. Finally, we provide a taxonomy for stream- and kernel-triggered abstractions, including disambiguation of similar semantic terms, and consider directions for future standardization in MPI-5.
@@ 1 Introduction
Most MPI implementations now include support for GPU data paths, enabling standard MPI host-based APIs to move data to and from GPU memory. This feature is generally termed GPU-aware MPI and significant progress has been made toward standardizing this implementation feature through work on MPI memory types [11]. Importantly, the MPI operation stages of initiation, starting, completion, and freeing [10] are all conducted by code running on the host CPU in this approach; GPU code that needs to cause or coordinate MPI operation state changes must do so indirectly by synchronizing with CPU code that makes those updates on behalf of the GPU code.
@@
The myriad proposed approaches for addressing these issues [1, 3, 5, 9, 10, 12, 13, 15, 16] have focused on GPU-triggered communication APIs in which host calls initialize and free MPI communication requests and the GPU control path initiates and in some cases completes MPI communication requests. When the GPU control path used for initiation and completion is the GPU stream scheduler, here and elsewhere this is termed stream-triggered communication and generally does not involve GPU-callable routines; instead the host API interacts with the GPU stream control program to schedule implicit GPU initiation and completion of communication operations. In contrast, using GPU kernel thread execution for initiation and completion is generally termed kernel-triggered communication and requires the addition of explicit GPU-callable APIs for initiating and completing communication.
@@
  • Summarizes the key features of and classifies eight different MPI stream- and kernel-triggering API proposals using this taxonomy and compares and contrasts key requirements and API design elements of these approaches (Section 4);
@@ 2 Representative GPU-Triggered Communication Abstractions
To illustrate the wide range of design choices and challenges in this area, we begin by providing information on three proposals for GPU-triggered MPI communication. Specifically, we discuss the stream triggered, two-sided communication support in recent MPICH implementations [16], the stream-triggered MPI one-sided communication support recently proposed by researchers at Hewlett Packard Enterprise (HPE) [13], and extensions to the MPI 4.0 Partitioned Communication interface [10] to support GPU kernel triggering of MPI communication [8]. We also present simple ping-pong communication for each proposal to further elucidate their proposed approaches. Together, this section highlights and provides essential background on the diverse design choices we explored in designing the GPU triggered communication taxonomy described below in Section 3. Note that while the examples in this section utilize CUDA APIs, appropriate ROCM APIs could be utilized instead.
@@ 2.1 MPICH Triggering
MPICH provides explicit enqueue variants of the supported GPU-triggered operations that return immediately to the host; GPU synchronization routines must be used to await for completion of these operations on the GPU. MPICH implementers considered making MPI_Send and MPI_Recv on communicators with GPU streams attached have enqueue semantics (similar to the previously-proposed MPI-GDS API [15], but in the end rejected this because it would change the semantics of well-known MPI APIs. Finally, MPICH stream communicators also include additional features to support APIs similar to the previous MPI Endpoints proposal [2].
@@ 23 MPIX_Send_enqueue(dst_buf, 1, MPI_INT, 0, 123,
  • MPICH’s stream communicators enqueued operations require the calling process must explicitly synchronize with the underlying GPU stream using GPU-specific operations to ensure completion of the corresponding MPI operation in the GPU execution context.
@@ 2.2 HPE One-Sided
The HPE one-sided GPU triggering API [13] supports stream-triggered, one-sided communication [^1] HPE also provides stream-triggered, two-sided communication, not considered here.. The stream abstractions included in HPE’s approach rely on GPU-NIC asynchronous stream triggering that integrates the NIC command queue and the GPU stream. In this approach, the CPU puts commands in the Cray Cassini NIC’s command queue and then GPU stream operations triggers these commands. This removes unnecessary synchronization on the CPU. HPE’s one-sided API is designed to allow for GPU-NIC progress where the host CPU is fully bypassed in progressing communication. HPE’s previous two-sided GPU triggering APIs [12] required host progress to receive messages due to the need for tag magtching and unexpected message handling.
@@ 20 MPIX_Win_complete_stream(win, stream);
  • The MPI_Win_complete_enqueue initiates operations initialized by the host in the command queue. This utilizes the GPU and NIC without weakening MPI RMA synchronization semantics, and in reality strengthens MPI RMA semantics by always deferring one-sided operations until MPI_Win_complete_ enqueue is executed by the GPU stream control program.
@@ 38 MPIX_Wait(&srequest, MPI_STATUS_IGNORE);
The MPI-ACX implementation [1] includes a prototype of the partitioned communication detailed in MPI-4. Partitioned point-to-point communication was specifically added to MPI-4 for a variety of reasons, including providing matching semantics (matching at initialization) and lightweight interfaces for data transfer initiation  [^2] It also removes the restrictions of persistent send and receive operations—MPI_Send_init and MPI_Recv_init—from MPI-1, which are forbidden to communicate and do not match only once. These fail to offer a channel semantic or obviate receive-queue matching. By contrast, single partitioned mode provides a uni-directional channel concept.. A specific nuance in this approach is that both MPIX_ Parrived and MPIX_Pready can be called from the host and the CUDA device. There are device-variants of the host requests to support the device-invoked operations.
Notably, the absence of a strong progress guarantee in MPI limits the guarantee that the partitioned communication initialization phase (i.e., MPI_Psend_init has already been matched proactively with MPI_Precv_init) when MPI_Start is called; it can instead occur at the first MPI_Wait for both sides of the partitioned channel. As such, proposals were offered for MPI-4.1 denoted MPI_Pbuf_prepare to guarantee initialization completion prior to first use and ensure readiness of buffers upon reuse [6]. This has not yet been adopted for the standard, but appears essential to avoid unexpected messages that would introduce complex logic that is considered untenable in GPU kernels.
@@
  • The host-based MPIX_Request object is used for regular MPI functions such as MPIX_Start and MPIX_Wait.
  • Different request types and funcation signatures are needed for GPU and CPU interfaces due to differences between the state needed in each context.
@@ 3 A Taxonomy of MPI GPU-Triggered Communication
Based on a careful examination of both the MPI standard and prior proposals for GPU-triggered communication, we created the following categories for classifying key characteristics of MPI GPU-triggered communication proposals in four different overarching areas:
@@
Does the API support the GPU stream control path or kernel control path for initiating and completing MPI communication operations? In the case of APIs where both stream and kernel triggering are supported, we split the API into separate stream- and kernel-triggered sub-APIs and classify them separately because of the frequent semantic and API differences between the such approaches in other categories.
@@
Does the GPU triggering API reuse existing MPI communication operations or abstractions to enable GPU triggering of communication operation initiation or completion, or does it instead create new MPI operations or abstractions?
@@
Separate MPI Operation Initialization and Initiation:
Does the API use a single call to execute the starting stages of MPI communication operations or does it separate this into explicit API calls to initialize and initiate MPI communication operations?
@@
Does the API specify operations enabling a single GPU communication operation to communicate with multiple processes? We classify collective support as either full, group, or none, cooresponding to (1) an API that supports the full semantics of MPI two-sided collecgtives, (2) an API that supports collective pairwise exchanges between subgroups of a communicator such as one-sided group or two-sided neighbor collectives, or (3) no specified support for collective communication.
@@
If the API includes an explicit sequencing abstraction, are communication operations sequenced using that abstraction fully ordered or partially ordered (e.g., using a fencing or graph abstraction)?
@@
Does the API enable message initiation and completion by the GPU and network device without the involvement of the host CPUS? This is generally accomplished by supporting one-sided data movement that avoids the control flow complexity of matching in the receive path.
@@ 4.1 API Classification
We focus on the eight proposed GPU-triggered MPI programming interface extensions listed below:
@@
– stream-triggered two-sided communication operations as implemented in the MPI-ACX GPU-triggering interface extension for NVIDIA network devices that support MLX4 device functionality.
@@
– a GPU triggering interface extension for the MPICH MPI implementation [7].
@@
Note that multiple variants of some of these designs have been proposed in the MPI Forum Working groups, BOFs, and other venues; we focus on the above listed API descriptions and implementations as representative of the main existing approaches at the time of this publication.
@@
Used APIs Semantics and Initiate Support
@@
MPICH Stream Yes No2 No Full No
@@ 4.2 API Feature Comparisons
Almost every API allows the GPU to check for or wait for the completion of every MPI operation that they initiate, allowing the GPU to avoid returning to the host execution context to ensure communication completion. The main exception is partitioned communication, where receives completion can be tested on the GPU using Parrived, but there is no equivalent call for testing for the completion of sends initiated from the kernel using Pready (the only available option is to check if all send partitions have completed using a single call to Wait or Test or one of their variants).
While several APIs propose collective support, only limited support for any GPU-triggered collective communication is available in the form of point-to-point group communication in the HPE one-sided triggering API. This is likely due to the complexity of combining complex collective communication with the constrained GPU and NIC control paths.
@@
Implementation availability also varies significantly. For instance, Delorean is just a specification with no known implementation; MPI-ACX is fully available and we were able to run the ping-ping example on representative hardware. Few implementations support full GPU/NIC offload of communication, with most APIs designed with the assumption that the host will be involved in data movement.
@@ 5 Key Gaps
Based on the classifications described in the previous section, this section highlights covers gaps and perceived limitations observed in prior work and prototypes as well as the MPI standard.
@@ 5.1 Limitations in the MPI Standard
In the MPI Standard, we observe that there are two main areas that needs to be addressed.
@@
    Problems surrounding persistence:
    The separation of initialization and initiation in GPU triggering APIs make persistent operations attractive for this purpose, as demonstrated by partitioned communication. Unfortunately, MPI-1 specified persistent send and receive operations cannot communicate and their initialization does not result in matching; indeed, persistent and non-persistent operations can match each time they are initiated.
@@
    Lack of MPI_Pbuf_prepare and similar operations for two-sided communication:
    There are two separate issues that have been conflated in the prior discussions of MPI_Pbuf_prepare and MPI_Pbuf_prepareall, yet to be standardized [6]:
@@
    The former evidently subsumes the latter; otherwise, matching/synchronization would have to be done prior to every call to MPI_Start, which is likely to be unacceptably high.
    Further, because of the allowance for weak progress in MPI implementations, there is no guarantee even in two sided operations that match at initialization (e.g. partitioned communication) that the initial match has occured between before the first complete communication. This means that initial communication can happens before any agreement on the readiness of the receive buffer; for instance, agreement and transmission could first happen during the first wait calls, or the initial transfers could be done differently than subsequent ones.
    In addition, the desire for these conditions is not unique to partitioned communication — the various stream-triggered communication proposals (MPICH, HPE two-sided, MPI-ACX enqueuing, etc.) would want this functionality to push their data movement through the one-sided path as well.

### v3

@@ Abstract
GPU-enhanced architectures are now dominant in HPC systems, but message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches that lower such costs. We compare and contrast stream/graph-, kernel-triggered, and GPU-initiated MPI communication abstractions, whose principal purpose is to enhance the performance of communication when GPU kernels create or consume data for transfer through MPI operations. Researchers and practitioners have proposed multiple potential APIs for GPU-involved communication that span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, and explicit MPI stream/queue objects. Designs breaking backward compatibility with MPI are duly noted. Some of these strengthen or weaken the semantics of MPI operations. A key contribution of this paper is to promote community convergence toward common abstractions for GPU-involved communication by highlighting the common and differing goals and contributions of existing abstractions. We describe the design space in which these abstractions reside, their implicit or explicit use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discuss their potential for added performance, how usable these abstractions are, and where functional and/or semantic gaps exist. Finally, we provide a taxonomy for these abstractions, including disambiguation of similar semantic terms, and consider directions for future standardization in MPI-5.
Keywords:
Stream-triggered message passing, Kernel-triggered message passing, GPU-initiated message passing, MPI, GPU, MPI+X, partitioned communication
@@ 1 Introduction
Most MPI implementations now include support for GPU data paths, enabling standard MPI host-based APIs to move data to and from GPU memory. This feature is generally termed GPU-aware MPI and significant progress has been made toward standardizing this implementation feature through work on MPI memory types [13]. Importantly, the MPI operation stages of initialization, starting, completion, and freeing [12] [^1] The term ‘initiating’ refers to a sequence of the initialization and starting phases. are all conducted by code running on the host CPU in this approach; GPU code that needs to cause or coordinate MPI operation state changes must do so indirectly by synchronizing with CPU code that makes those updates on behalf of the GPU code.
@@
The myriad proposed approaches for addressing these issues [1, 3, 5, 11, 12, 14, 15, 17, 18] have focused on GPU-triggered communication APIs in which host calls initialize and free MPI communication requests and the GPU control path starts and in some cases completes MPI communication requests. When the GPU control path used for starting and completion is the GPU stream scheduler, here and elsewhere this is termed stream-triggered communication and generally does not involve GPU-callable routines; instead the host API interacts with the GPU stream control program to schedule implicit GPU starting and completion of communication operations. In contrast, using GPU kernel thread execution for starting and completion is generally termed kernel-triggered communication and requires the addition of explicit GPU-callable APIs for starting and completing communication. As opposed to these options, Intel’s approach includes GPU-initiated APIs (in which all parameters to the operation are provided in the GPU) and GPU-synchronization APIs [9, 10].
@@
  • Summarizes the key features of and classifies nine different MPI stream- and kernel-triggering API proposals using this taxonomy and compares and contrasts key requirements and API design elements of these approaches (Section 4);
@@ 2 Representative GPU-Triggered Communication Abstractions
To illustrate the wide range of design choices and challenges in this area, we begin by providing information on four proposals for GPU-triggered MPI communication. Specifically, we discuss the stream triggered, two-sided communication support in recent MPICH implementations [18], the stream-triggered MPI one-sided communication support recently proposed by researchers at Hewlett Packard Enterprise (HPE) [15], extensions to the MPI 4.0 Partitioned Communication interface [12] to support GPU kernel triggering of MPI communication [8], as well as the GPU-initiated model proposed by Intel [9, 10].
We also present simple ping-pong communication for each proposal to further elucidate their proposed approaches. Together, this section highlights and provides essential background on the diverse design choices we explored in designing the GPU triggered communication taxonomy described below in Section 3.
@@ 2.1 MPICH Triggering
MPICH provides explicit enqueue variants of the supported GPU-triggered operations that return immediately to the host; GPU synchronization routines must be used to await for completion of these operations on the GPU. MPICH implementers considered making MPI_Send and MPI_Recv on communicators with GPU streams attached have enqueue semantics (similar to the previously proposed MPI-GDS API [17]), but in the end rejected this because it would change the semantics of well-known MPI APIs. Finally, MPICH stream communicators also include additional features to support APIs similar to the previous MPI Endpoints proposal [2].
@@ 23 MPIX_Send_enqueue(dst_buf, 1, MPI_INT, 0, 123,
  • MPICH’s stream communicators enqueued operations require that the calling process must explicitly synchronize with the underlying GPU stream using GPU-specific operations to ensure completion of the corresponding MPI operation in the GPU execution context.
@@ 2.2 HPE One-Sided
The HPE one-sided GPU triggering API [15] supports stream-triggered, one-sided communication [^2] HPE also provides stream-triggered, two-sided communication, not considered here.. The stream abstractions included in HPE’s approach rely on GPU-NIC asynchronous stream triggering that integrates the NIC command queue and the GPU stream. In this approach, the CPU puts commands in the Cray Cassini NIC’s command queue and then GPU stream operations triggers these commands. This removes unnecessary synchronization on the CPU. HPE’s one-sided API is designed to allow for GPU-NIC progress where the host CPU is fully bypassed in progressing communication. HPE’s previous two-sided GPU triggering APIs [14] required host progress to receive messages because of the need for tag matching and unexpected message handling.
@@ 20 MPIX_Win_complete_stream(win, stream);
  • The MPI_Win_complete_stream starts communication operations in the
    command queue initialized by the host (e.g., using MPI_Put). This utilizes the GPU and NIC without weakening MPI RMA synchronization semantics, and in reality requires stronger MPI RMA synchronization semantics by always deferring one-sided operations until MPI_Win_complete_ stream is executed by the GPU stream control program. That is, in the HPE one-sided API, the communications indicated by MPI_Put calls cannot legally occur until MPI_Win_complete_stream is executed in stream order by the GPU control program.
@@ 38 MPIX_Wait(&srequest, MPI_STATUS_IGNORE);
The MPI-ACX implementation [1] includes a prototype of the partitioned communication detailed in MPI-4. Partitioned point-to-point communication was specifically added to MPI-4 for a variety of reasons, including providing matching semantics (matching at initialization) and lightweight interfaces for data transfer initiation [^3] It also removes the restrictions of persistent send and receive operations—MPI_Send_init and MPI_Recv_init—from MPI-1, which are forbidden to communicate and do not match only once. These fail to offer a channel semantic or obviate receive-queue matching. By way of contrast, single partitioned mode provides a uni-directional channel abstraction.. A specific nuance in this approach is that both MPIX_ Parrived and MPIX_Pready can be called from the host and the CUDA device. There are device-variants of the host requests to support the device-invoked operations.
Notably, the completion of an MPI_Psend_init call does not guarantee that a receiver has been matched prior to calling MPI_Start; it can instead occur at the first MPI_Wait for both sides of the partitioned channel. As such, proposals were offered for MPI-4.1 denoted MPI_Pbuf_prepare to guarantee initialization completion prior to first use and ensure readiness of buffers upon reuse [6]. This has not yet been adopted for the standard, but appears essential to avoid unexpected messages that would introduce complex logic that is considered untenable in GPU kernels.
@@
  • The host-based MPI request object is used for regular MPI functions such as MPIX_Start and MPIX_Wait.
  • Different request types and function signatures are needed for GPU and CPU interfaces due to differences between the state needed in each context.
2.4 Intel GPU-Initiated Communication
Intel provides a GPU-initiated communication feature [10] that appears similar to HPE’s one-sided communication prototype. Intel’s additions to its MPI implementation [9] is kernel-triggered, and uses existing SYCL queues for communication ordering instead of an explicit MPI ordering object (as MPICH does). Figure 4 demonstrates the ping-pong exemplar following Intel’s approach that reduces the number of MPI functions to only MPIX_Put_notify and MPIX_- Win_get_notify instead putting work on the GPU using a SYCL queue object and adding them as events [^4] Personal Communication, Dr. Daniel J. Holmes, Intel Corp., June 14, 2024.
1void do_pingpong_mpi(int target, sycl::queue &q, [[maybe_unused]] void **ptrs, int threads=1) {
2 double latency = 0.0;
3 int tgt = target;
4 MPI_Win_fence(0, win);
5 for (int i=0; i<2;++i) {
6 int iter = iters[i];
7 if (rank == 0) {
8 sycl::event e = q.submit([&](sycl::handler &h) {
9 h.parallel_for_work_group(sycl::range(1), sycl::range(threads), [=](sycl::group<1> grp) {
10 grp.parallel_for_work_item([&](sycl::h_item<1> item) {
11 int id = item.get_global_id(0);
12 int v = 0; MPI_Count c = 0;
13 for (int i=1; i<iter; ++i) {
14 MPIX_Put_notify(&v, 1, MPI_INT, tgt, id, 1, MPI_INT, id, win);
15 while (c<i)
16 MPIX_Win_get_notify(win, id, &c);
17 } }); }); });
18 q.wait();
19 } else if (rank == target) {
20 q.submit([&](sycl::handler &h) {
21 h.parallel_for_work_group(sycl::range(1), sycl::range(threads), [=](sycl::group<1> grp) {
22 grp.parallel_for_work_item([&](sycl::h_item<1> item) {
23 int id = item.get_global_id(0);
24 int v = 0; MPI_Count c = 0;
25 for (int i=1; i<iter; ++i) {
26 while (c<i)
27 MPIX_Win_get_notify(win, id, &c);
28 MPIX_Put_notify(&v, 1, MPI_INT, 0, id, 1, MPI_INT, id, win);
29 } }); }); }).wait();
30 } }
31 MPI_Win_fence(0, win);
32 for (i = 0; i < threads; ++i)
33 MPIX_Win_set_notify(win, i, (MPI_Count)0);
34}
Figure 4: Ping Pong using Intel GPU-Initiated MPI [9, 10]
@@ 3 A Taxonomy of MPI GPU-Triggered Communication
Based on a careful examination of both the MPI standard and prior proposals for GPU-triggered communication, we defined the following categories for classifying key characteristics of MPI GPU-triggered communication proposals in four different overarching areas:
@@
Does the API support the GPU stream control path or kernel control path for starting and completing MPI communication operations? In the case of APIs where both stream- and kernel-triggering are supported, we split the API into separate stream- and kernel-triggered sub-APIs and classify them separately because of the frequent semantic and API differences between the such approaches in other categories.
@@
Does the GPU-triggering
API reuse existing MPI communication operations or abstractions to enable GPU triggering of communication operation starting or completion, or does it instead create new MPI operations or abstractions?
@@
Separate MPI Operation Initialization and Starting:
Does the API use a single call to execute the initiate stages of MPI communication operations or does it separate this into explicit API calls to initialize and start MPI communication operations?
@@
Does the API specify operations enabling a single GPU communication operation to communicate with multiple processes? We classify collective support as either full, partial, group, or none, corresponding to (1) an API that supports the full semantics of MPI two-sided collective operations, (2) an API that supports a subset of the MPI two-sided collective operations, (3) an API that supports collective pairwise exchanges between subgroups of a communicator such as one-sided group or two-sided neighbor collectives, or (4) no specified support for collective communication.
@@
If the API includes an explicit sequencing abstraction, are communication operations sequenced using it fully ordered or partially ordered (e.g., using a fencing or graph abstraction)?
@@
Does the API enable message initiation or starting and completion by the GPU and network device without the involvement of the host CPUs? This is generally accomplished by supporting one-sided data movement that avoids the control flow complexity of matching in the receive path.
@@ 4.1 API Classification
We focus on the nine proposed GPU-triggered MPI programming interface extensions listed below:
@@
– stream-triggered, two-sided communication operations as implemented in the MPI-ACX GPU-triggering interface extension for NVIDIA network devices that support MLX4 device functionality.
@@
– a GPU-triggering interface extension for the
MPICH MPI implementation [7].
@@
Intel GPU-Initiated [9, 10]
– a GPU-initiated extension for Intel MPI [^5] Personal Communication, Dr. Daniel J. Holmes, Intel Corp., June 14, 2024.
Note that while multiple variants of some of these designs have been proposed in the MPI Forum Working groups, BOFs, and other venues, we focus on the above listed API descriptions and implementations as representative of the main existing approaches at the time of this publication.
@@
Used APIs Semantics and Start Support
@@
MPICH Stream Yes No2 No Full Partial
@@ 4.2 API Feature Comparisons
Almost every API allows the GPU to check for or wait for the completion of every MPI operation that they either initiate or start, enabling the GPU to avoid returning to the host execution context to ensure communication completion. The main exception is partitioned communication, where receives completion can be tested on the GPU using Parrived, but there is no equivalent call for testing for the completion of sends initiated from the kernel using Pready (the only available option is to check if all send partitions have completed using a single call to Wait or Test or one of their variants).
While several APIs propose collective support, only limited support for any GPU-triggered collectives is available, e.g., MPIX_Allreduce_enqueue in MPICH and point-to-point group communication in the HPE one-sided triggering API. This is likely due to the complexity of combining complex collective communication with the constrained GPU and NIC control paths.
@@
Implementation availability varies significantly. For instance, Delorean is just a specification with no known implementation; Intel’s GPU support is currently in “preview,” while MPI-ACX is fully available and so we were able to run the ping-ping example on representative hardware. Few implementations support full GPU/NIC offload of communication; most APIs designed with the assumption that the host will be involved in data movement.
@@ 5 Key Gaps
Based on the classifications described in the previous section, this section covers gaps and perceived limitations observed in prior work and prototypes as well as the MPI standard.
@@ 5.1 Limitations in the MPI Standard
In the MPI Standard, there are two main areas that needs to be addressed.
@@
    Problems surrounding persistence: The separation of initialization and starting in GPU triggering APIs make persistent operations attractive for this purpose, as demonstrated by partitioned communication. Unfortunately, MPI-1 specified persistent send and receive operations cannot communicate at initialization nor does their initialization result in matching; indeed, persistent and non-persistent operations can match each time they are started.
@@
    Lack of MPI_Pbuf_prepare* operations for two-sided communication: Two separate issues that have been conflated in prior discussions of MPI_Pbuf_prepare and MPI_Pbuf_prepareall, yet to be standardized [6]:
@@
    The former evidently subsumes the latter; otherwise, matching/synchronization would have to be done prior to every call to MPI_Start, which is likely to be unacceptably costly. Further, because of the allowance for weak progress in MPI implementations, there is no guarantee even in two-sided operations that match at initialization (e.g., in partitioned communication) that the initial match has occurred between before the first complete communication. Thus, initial communication can occur before any agreement on the readiness of the receive buffer (e.g., agreement and transmission could first transpire during the first wait calls, and/or the initial transfers could be orchestrated differently than subsequent ones).
    In addition, the desire for these conditions is not unique to partitioned communication—the various stream-triggered communication proposals
    (MPICH, HPE two-sided, MPI-ACX enqueuing, etc.) would want this functionality to push their data movement through the one-sided path as well.